# Stuffed Mamba: State Collapse and State Capacity of RNN-Based Long-Context Modeling

### v1

@@ 4.2 What is the Cause of State Collapse?
(c) The statistics of $x_{t}$ over time. Each curve represents a head.
@@
Figure 6: The value of the components in the update rule ($\Delta_{t}$, $B_{t}$, and $x_{t}$) on some heads with SC in the 38th layer in Mamba-2 370M. The red dotted line indicates the training length.
4.2.1 State Collapse as a Result of State Overparameterization
Figure 7: The cumulative decay of the first token ($\alpha_{1:t}$) over time. Each line represents a head. The statistics are collected from the 38th layer of Mamba-2 370M.
@@ 4.2.1 State Collapse as a Result of State Overparameterization
Figure 8: The passkey retrieval results of intermediate checkpoints during the pre-training of Mamba-2 370M on 512 sequence length. SC only occurs in the model beyond a certain amount of training data.
Here, we present a high-level explanation for SC. We argue that SC arises from state overparameterization relative to the training length. In other words, the state capacity is excessively large for the training length, allowing the model to achieve strong language modeling performance without learning how to forget when the state is about to overflow. To support this argument, we formulate the hidden state as a weighted sum of previously inserted information:
Therefore, $\alpha_{i:t}$ describes the memory strength about the $i$-th token at $t$ time step. Figure 8 shows the memory strength of the first token at different time steps, and we find that the exploded heads (heads 2, 4, and 7 in the 38th layer) have a strong inclination toward retaining all information within the training length, with a memory strength of over 0.8 at $t$=8K. This implies that the model has not learned to forget information (by producing a smaller decay $\alpha_{j}$) in a manner to avoid overflowing the state with too much information. Furthermore, we pre-train Mamba-2 from scratch with $T_{\text{train}}=512$ and evaluate the intermediate checkpoints on passkey retrieval, as reported in Figure 8. It shows that SC is only exhibited by checkpoints beyond a certain amount of training, which coincides with behaviors of overfitting—a result of overparameterization. One can also notice that the overfitted checkpoint outperforms earlier checkpoints on shorter sequences, which further strengthens the hypothesis that the model converges to less forgetting. Finally, as we will show in Section 4.3.2, for any given training length, there exists a state size where SC will be exhibited if and only if the model’s state size is greater.
4.3 How to Mitigate State Collapse?
Based on the analyses in the previous section, we propose several SC mitigation methods to make the model generalize better along the sequence length. In brief, we propose three training-free methods by modifying the update rule to avoid overflowing the state. Additionally, we directly train on longer sequences to encourage learning to smoothly forget the earliest information when the context is too long.
4.3.1 Training-Free Mitigation Methods
Method 1: Forget More and Remember Less
Since the variance of the state explodes during SC, we can reduce this by increasing the amount of state decay (i.e., forget more) or reducing the amount of inserted information (i.e., remember less). Based on the analysis in Section 4.2, we choose to intervene in the components $B_{t}$ and $\alpha_{t}$, which control the insertion strength and memory decay strength, respectively. Existing works have experimented with modifying $\Delta_{t}$, but it controls both the insertion and decay strength, making it hard to analyze and control.
Method 2: State Normalization
The main idea is to normalize the state after each update to ensure that the state’s norm is always below a threshold $p\in\mathbb{R}$. Specifically, we decay the state $h_{t}$ at each time step to ensure that $\|h_{t}\|\leq p$. Thus, we get the following update rule.
$\hat{h}_{t}$ $=h_{t-1}\overline{A}_{t}+\overline{B}^{T}x_{t}$ (12)
$h_{t}$ =$=\begin{cases}\hat{h}_{t}p/\|\hat{h}_{t}\|&\text{ if }\|\hat{h}_{t}\|>p\\ \hat{h}_{t}&\text{ if }\|\hat{h}_{t}\|\leq p\\ \end{cases}$ (13)
It is worth mentioning that this converts the model into a non-linear RNN and cannot be parallelized in the same manner as the original model, making it much slower at pre-filling.
Method 3: Sliding Window by State Difference
$h_{t}^{(r)}=\sum_{i=t-r+1}^{t}\alpha_{i:t}\hat{R}_{i}=\sum_{i=1}^{t}\alpha_{i:t}\overline{B}_{i}^{T}x_{i}-\alpha_{t-r+1:t}\sum_{i=1}^{t-r}\alpha_{i:t-r}\overline{B}_{i}^{T}x_{i}=h_{t}-\alpha_{t-r+1:t}h_{t-r}$ (14)
During streaming generation, we only have to maintain $(h_{t-1},h_{t-r},\alpha_{t-r+1:t})$ [^3] We also have to cache the last $r$ token IDs, but their size is negligible compared to $h_{t-1}$ and $h_{t-r}$., and advance each of them in parallel. However, directly computing $\alpha_{t:t-r}$ may suffer from instability due to floating-point imprecision. Therefore, we maintain $\Delta_{t-r:t}=\sum_{i=t-r}^{t}\Delta_{t}$ instead, and re-compute $\alpha_{t-r:t}=\exp\left(-\Delta_{t-r:t}\exp(A)\right)$ at every step, which incurs minimal computational cost.
This method applies to all RNNs that can be written as a weighted sum, which includes RWKV 5 and 6, RetNet, GLA, etc. It doubles the computation and memory cost for generation, but we believe that it is an acceptable trade-off because RNNs have a very low generation cost compared to transformer-based models and the context processing cost is unchanged.
4.3.2 Training on Longer Sequences
Based on the hypothesis that SC is caused by state overparameterization (described in Section 4.2.1), we can simply train on lengths that exceed the state capacity, which we conduct in this section.
Data Engineering
To ensure that the data contains as much long-term structure as possible, we filter out sequences with less than 4K tokens. Buckman & Gelada (2024) have shown that this is critical for training effective long-context models. Although we train on sequences longer than 4K tokens, we do not use a higher length threshold because the above threshold already removes about 97.6% of the data in the original corpus. To train on longer sequences, we simply concatenate sequences and delimit them with a special EOS (End-of-Sequence) token.
Truncated Backpropagation Through Time
In the vanilla Mamba-2, the states are initialized to zeros for each data sample. Instead, we initialize the states as the final state of the previous sequence. This is equivalent to concatenating multiple sequences, but stopping the backpropagation of gradients at certain intervals. This technique has been shown to help extend the context length of RNNs (Yang et al., 2023) and alleviate the memory cost of caching activations for computing gradients. Based on Yang et al. (2023) and our preliminary tests, we use concatenate 12 sequences with this technique by default.
5 State Capacity
Based on the discussion in Section 4.2.1, SC occurs if and only if the training length contains less information than the state capacity. Thus, we can indirectly estimate state capacity by sweeping different training lengths for different state sizes, and observing when the state does not collapse. In this section, we empirically investigate the relationship between state capacity and state size.
Specifically, we conduct the same training as in Section 4.3.2. To determine whether a state has collapsed, we feed the “newlines” prompt with 1M tokens to the model, and define collapse as the point where perplexity is more than 2x the maximum perplexity in its training length. We train multiple Mamba-2 with different state sizes and training lengths, and we regard the minimum training length at which SC does not occur as the state capacity.
5.1 State Capacity in Passkey Retrieval
@@
The language modeling performance may not reflect the downstream capabilities well (Fu et al., 2024). Therefore, we also search for the state capacity on the passkey retrieval task. Similar to the previous section, we train with different lengths for different state sizes and identify the maximum context length where the model has an accuracy over 95%, which we regard as the state capacity in passkey retrieval. In this task, the noisy context is repetitive, thus, the amount of contextual information is largely independent of the context length, therefore, the capacity should grow roughly exponentially with the state size.
It is worth emphasizing that, if we train Mamba-2 on passkey retrieval data, the model can theoretically handle infinitely long contexts by ignoring all irrelevant tokens. Here, the model is only trained with the next token prediction objective, which means the model will not ignore the irrelevant context, and the ability to retain information for extended time emerges from language modeling.
@@ 6 Experiments
We briefly describe the experimental details for length extrapolation experiments.
@@
We start from RedPajama-V2 (Computer, 2023), an open dataset with 30T tokens extracted from CommonCrawl [^4] https://commoncrawl.org/, and we perform deduplication to ensure data quality. During evaluation, we sample documents longer than 16K tokens and concatenate them if it is not long enough.
@@
We experiment with seven model configurations with different state sizes to find the relationship between state capacity and size. For each of them, we perform an extensive search with training lengths up to 256K tokens. To save cost, we continue pre-train from three official checkpoints of Mamba-2 of size 130M, 370M, and 780M. They are all trained on 8K sequences. The other three model configurations (36M, 47M, and 85M) are trained from scratch. The detailed configurations are given in Appendix F.1.
Hyperparameters
We use the WSD (Hu et al., 2024) with 10% decay steps. This scheduler is chosen because it is competitive with the commonly used cosine scheduler while allowing simple resumption from intermediate checkpoints, saving large amounts of computational resources. We report the result of the best checkpoint selection by validation on passkey retrieval. More hyperparameters are reported in the Appendix F.
7 Results
7.1 Training-Free Length Generalization
Figure 9 reports the result of the training-free length generalization methods on Mamba-2 780M. We can see that while LongMamba [^5] https://github.com/jzhang38/LongMamba can greatly improve the length generalizability of the model by more than 3x, it causes noticeably greater perplexity on shorter sequences and still inevitably exhibits SC. All our methods successfully suppress SC, allowing the model to generalize to more than 64K tokens, although State Normalization greatly underperforms other methods on shorter sequences. One explanation for this underperformance is that normalizing collapsing states changes the ratio of norms between heads, and this disrupts the learned mechanisms.
Figure 9: Result of training-free length generalization methods described in Section 4.3.1. Perplexity is computed on long documents from RedPajama (Computer, 2023). The red dotted line represents the training length.
7.2 Length Generalization by Training on Longer Sequences
(a) 130M.
(b) 370M.
Figure 10: The perplexity as a function of token position for Mamba-2 130M on the “newlines” prompt with three different training lengths.
Figure 11: Passkey retrieval Performance of the best checkpoint of Mamba-2 370M. The x-axis is in the log scale.
@@ 7.2 Length Generalization by Training on Longer Sequences
In Figure 11, we plot the language modeling perplexity as a function of token position for Mamba-2 130M and 370M with different training lengths. We can see that for each model size, there is a training length threshold, beyond which the model has much better length extrapolation, which supports our arguments discussed in Section 4.3.2.
7.3 State Capacity as a Function of State Size
Figure 12: The left figure shows the training length beyond which the model does not exhibit SC. The right figure shows the memory capacity on passkey retrieval as a function of state size.
@@ 7.3 State Capacity as a Function of State Size
Figure 12 shows the state capacity of Mamba-2 on language modeling and passkey retrieval. The rightmost data point in both plots corresponds to Mamba-2 370M. We have confirmed that 780M also exhibits SC at training lengths below 128K, but do not have enough resources to train the model beyond this length. The results establish a linear relationship $T_{\text{train}}=5.172\cdot S-4.469$ between the length $T_{\text{train}}$ at which SC stops occurring and the state size $S$.
@@
The second plot of Figure 12 shows that the capacity of Mamba-2 on passkey retrieval is exponential concerning the state size. This is because the amount of information in the context does not increase with its length. In other words, we are storing a constant amount of information while the number of combinations of the state grows exponentially with the number of elements. Figure 11 shows the best checkpoint of Mamba-2 370M on passkey retrieval. The result is very promising because, to the best of our knowledge, no previous models with less than 1B model parameters have near-perfect accuracy at 128K tokens in this task.
@@ 8 Conclusion
This paper discovers and presents the first systemic study on state collapse (SC), a phenomenon in RNNs that causes length generalization failure. With an inspection of the activations and controlled experiments on Mamba-2, we conclude that this phenomenon is caused by an overparameterized state with excessive state capacity. Based on the analyses, we have proposed three training-free methods for mitigating SC up to 1M tokens. Then, we show that SC can be mitigated by training on context lengths that exceed the state capacity. With this insight, we empirically estimate the state capacity of Mamba-2 on language modeling and the passkey retrieval task. With some simple data engineering and state initialization tricks, we achieve much better performance with Mamba-2 on the passkey retrieval task than existing models. Our results indicate that Mamba-2 not only is highly efficient in handling long sequences but also has great performance potential.
Limitations
All models studied in this work can be seen as a specific case of linear attention models, whose recurrent state is decayed by an element-wise or scalar gate (they can be viewed as variants of Gated Linear Attention (Yang et al., 2023)). We have chosen to study these models because of their strong capabilities, yet, some conclusions may not be directly transferred to other variants of RNNs.
Our continued pre-training approach for extending the context length of RNNs is rather expensive, some of the models require training with up to 50B tokens, which is 1/6 of their pre-training amount of data. Also, to ensure simplicity, controllability, and generality, we have not used more advanced techniques for training long-context models, such as upsampling longer data samples, better data order or format, using data with more long-distance dependencies, etc.
The passkey retrieval task that we used extensively in this study is very simple. Hence, high accuracy on this task may not reflect the capabilities of the model on real-world long-context tasks, because that requires more advanced capabilities such as high-resolution retrieval, reasoning, state-tracking, etc. The result is nonetheless promising because it indicates that the model can recall the correct information and further capabilities may be achieved by building on the recalled information.
SC is somewhat prompt-dependent. While we found that for models with greatly overparameterized states, SC is highly consistent across different prompts, certain models with less overparameterization may exhibit SC on some prompts while successfully extrapolating indefinitely on others. Attributing SC to specific features of the prompt is a promising future research direction.

### v2

@@ 3.1 The Importance of Forgetting
Figure 3: The retention strength of the first token ($\alpha_{1:t}$) over time. Each curve represents a head.
Figure 4: LM loss of Mamba-2 370M at different positions when inducing forgetting (see Section 3.2.2).
Figure 5: The mean and variance of the first 8 heads in layer 38 of Mamba-2 370M. It exhibits a clear explosion when $t$ is greater than the training length.
3.2 Evidence for the Inability to Forget
Here, we provide empirical evidence that confirms that Mamba-2 has failed to learn how to forget past information.
3.2.1 Evidence 1: High Retention of the First Token
Based on Eq. 6, we can view $\alpha_{i:t}$ as the memory strength of the $i$-th token at $t$ time step. The retention strengths of earlier tokens are always smaller than those of more recent tokens. We find that some heads have a very strong inclination toward retaining all information within the training length. As an example, Figure 5 shows the cumulative decay of the first token in the first eight heads of the 38th layer, and three of the heads have a memory retention strength over 0.997333This is a cumulative product, so the decay at each time step is even closer to 1. at $t$=8K. Similar observations can be found in other heads and in other layers as well. This implies that the model has not learned to forget information (by producing a smaller $\alpha_{j}$), but it still has decent language modeling capabilities because the information of 8K tokens is typically not enough to overload the memory.
3.2.2 Evidence 2: Inducing Forgetting Can Improve Length Generalization
Here, we demonstrate that artificially inducing more forgetting without training can improve performance by reducing past memory interference.
Reduced Memory Retention and Insertion (RRI)
This method assumes that $\alpha_{t}$ and $B_{t}$ control the memory retention and insertion strength, respectively. We scale them with a multiplier smaller than 1. The actual multipliers used are $0.9999$ for $\alpha_{t}$ and $0.75$ for $B_{t}$, which is chosen by validation using average loss on pre-training data with 32K context length.
Sliding Window
$h_{t}^{(r)}=\sum_{i=t-r+1}^{t}\alpha_{i:t}\overline{B}_{i}x_{i}=\sum_{i=1}^{t}\alpha_{i:t}\overline{B}_{i}x_{i}-\alpha_{t-r+1:t}\sum_{i=1}^{t-r}\alpha_{i:t-r}\overline{B}_{i}x_{i}=h_{t}-\alpha_{t-r+1:t}h_{t-r}$ (8)
During streaming generation, we only have to maintain $(h_{t-1},h_{t-r},\alpha_{t-r+1:t})$ [^4] We also have to cache the last $r$ token IDs, but their size is negligible compared to $h_{t-1}$ and $h_{t-r}$, and advance each of them in parallel. However, directly computing $\alpha_{t:t-r}$ may suffer from instability due to floating-point imprecision. Therefore, we maintain $\Delta_{t-r:t}=\sum_{i=t-r}^{t}\Delta_{t}$ instead, and re-compute $\alpha_{t-r:t}=\exp\left(-\Delta_{t-r:t}\exp(A)\right)$ at every step, which incurs minimal computational cost. This method can be applied to all RNNs that can be written as a weighted sum.
Result
From Figure 5, one can see that the original model has the worst length generalization abilities. LongMamba (a length extrapolation method) (Zhang, 2023) and our methods for inducing forgetting can alleviate this generalization failure, although LongMamba and RRI compromise short-context performance due to weaker memory insertion. This further confirms the fact that memory over-retention is the culprit of this degradation.
3.3 The Manifestation of Over-Retention
[…]
Figure 6: The distribution of the channels in an exploding state (head 2 in layer 38) at two different time steps.
(a) $\Delta_{t}$ over time. Each curve represents a head.
(b) $B_{t}$ over time. Each curve represents a channel.
(c) $x_{t}$ over time. Each curve represents a head.
@@ 3.3 The Manifestation of Over-Retention
Figure 7: The value of various components in the update rule ($\Delta_{t}$, $B_{t}$, and $x_{t}$) on some heads with large retention values in the 38th layer in Mamba-2 370M. The red dotted line indicates the training length.
4 State Overparameterization
Here, we present a high-level explanation for why Mamba-2 over-retains memories: state overparameterization. The state is excessively large for the training length, allowing the model to achieve strong language modeling performance without learning how to forget when the state is overloaded with memories.
To support this hypothesis, we present two pieces of evidence: (1) Mamba-2 starts with the ability to forget, but slowly loses this ability as the amount of training data increases, which coincides with behaviors of overfitting, and (2) for any state size, there is a training context length threshold $T_{\text{forget}}$ such that Mamba-2 learns to forget if and only if $T_{\text{train}}>T_{\text{forget}}$, where $T_{\text{train}}$ denotes the training context length.
4.1 Evidence 1: More Training $\Rightarrow$ Less Forgetting
We pre-train Mamba-2 370M from scratch with $T_{\text{train}}=512$ using the RedPajama (Computer, 2023) corpus and evaluate the intermediate checkpoints on passkey retrieval, as reported in Figure 9. It shows that the model’s retrieval accuracy for contexts longer than the training length slowly decreases as we increase the amount of training data. Meanwhile, the model’s accuracy for contexts shorter than the training length increases. This indicates that the model converges toward more retention and less forgetting. Since language modeling loss is only computed for tokens within the training length, this behavior is induced by minimizing loss. This reduced forgetting leads to conflicts between token representations, impairing memory recall accuracy when an excessive number of tokens are inserted.
This behavior can be viewed as a kind of overfitting because the state’s distribution with short contexts is not varied enough for the model to generalize to the distribution in longer contexts. In other words, the state has too many parameters for the given training length.
@@ 4.1 Evidence 1: More Training $\Rightarrow$ Less Forgetting
Figure 8: Passkey retrieval results of intermediate checkpoints during the pre-training of Mamba-2 370M on 512 sequence length. Generalization failure only occurs in the model beyond a certain amount of training data.
Figure 9: The maximum context length of accurate passkey retrieval (i.e., $T_{\text{recall}}$ in Section 4.3) as a function of state size.
4.2 Evidence 2: Forgetting is Learned $\Leftrightarrow$ Sufficient Training Length
We empirically confirm that larger recurrent states require longer training contexts for the model to learn to forget. This is because the model will only learn to forget when the amount of contextual information exceeds the state capacity. This hypothesis implies the following law:
Let $N_{S}$ and $T_{\text{train}}$ denote the recurrent state size and training length, respectively, there exists a threshold $T_{\text{forget}}(N_{S})$ such that the model learns to forget if and only if $T_{\text{train}}>T_{\text{forget}}$.
We empirically validate this by sweeping different training lengths for different state sizes, and checking whether the model has successfully learned how to forget. Concretely, we train multiple Mamba-2 with different state sizes and training lengths to find the relationship between $T_{\text{forget}}$ and $N_{S}$. To determine whether the model has learned robust forgetting, we feed prompts with 1M tokens to the model and check if the model’s loss exceeds 2×2$\times$2 × the maximum loss within $T_{\text{train}}$ tokens at any point. The loss is averaged over 128 prompts. The result is reported in Section 5.1.
4.3 Maximum Recall Context Length
@@
The fact that the amount of information in $T_{\text{forget}}$ tokens exceeds the state’s capacity, does not necessarily imply that the model fails to recall information beyond the last $T_{\text{forget}}$ tokens, especially when there is a clear distinction between the target information and other contextual information. Therefore, we also search for the maximum context length from which the model can accurately perform passkey retrieval. We refer to this context length as the model’s maximum recall context length, denoted with $T_{\text{recall}}$. Similar to the previous section, we train with different lengths for different state sizes and identify the maximum context length where the model has an accuracy over 95% as $T_{\text{recall}}$. In this task, the noisy context is repetitive, thus, the amount of contextual information is largely independent of the context length. Therefore, ideally, the recall threshold should grow roughly exponentially with the state size. [^6] If we train Mamba-2 on passkey retrieval data, the model can theoretically handle infinitely long contexts. Here, the model is only trained with the next token prediction objective, which means the model will not ignore the irrelevant context, and the ability to retain information for extended time emerges from language modeling.
@@ 5 Experiments
We briefly describe the data and model configurations used to identify the forget threshold $T_{\text{forget}}$ and maximum recall context length $T_{\text{recall}}$. Due to limited space, more comprehensive experimental details are reported in Appendix F.
@@
We start from RedPajama-V2 (Computer, 2023), an open dataset with 30T tokens from the Internet, and perform deduplication to ensure data quality and discard documents that are too short.
@@
We experiment with six model sizes to find the relationship between state capacity and size. For each of them, we perform an extensive search with training lengths up to 256K tokens. To save cost, we continue pre-training from three official checkpoints of Mamba-2 (130M, 370M, and 780M). They were pre-trained with 8K sequences. The other model configurations (36M, 47M, and 85M) are trained from scratch.
(a) Mamba-2 130M
(b) Mamba-2 370M
Figure 10: LM loss at each token position for different training lengths. Evaluated on RedPajama.
Figure 11: Minimum training length at which the model learns robust forgetting (i.e., $T_{\text{forget}}$ in Section 4.2).
5.1 The Existence of Forget Threshold
@@
In Figure 11, we plot the language modeling perplexity as a function of token position for Mamba-2 130M and 370M with different training lengths. We can see that for each model size, there is a training length threshold, beyond which the model has much better length extrapolation, which supports our arguments discussed in Section 4.2.
5.2 Forget Threshold as a Function of the State Size
@@
Figure 11 shows the minimum training length needed for Mamba-2 to learn forgetting. The rightmost data point in the plot corresponds to Mamba-2 370M. We have confirmed that the 780M model (with a state size of 19.3M) also has poor length generalization at training lengths below 128K, but do not have enough resources to train the model beyond this length. The results establish a linear relationship $T_{\text{forget}}=5.172\cdot N_{S}-4.469$ between the length $T_{\text{train}}=T_{\text{forget}}$ at which the model can learn robust forgetting and the state size $N_{S}$. The $R^{2}$ value is over $0.999$. This indicates that to train a Mamba-2 with robust length generalization, one should use training lengths that grow linearly with the state size.
5.3 Maximum Recall Context Length as a Function of the State Size
@@
The second plot of Figure 9 shows the recall threshold of Mamba-2 in passkey. The maximum contexts length in which Mamba-2 can accurately retrieve 5-digit passkeys is exponential concerning the state size, the function is $T_{\text{recall}}=4.756\cdot(1.365^{N_{S}}-1)-0.742$, with an $R^{2}$ value over $0.999$. This is because the amount of information in the context does not increase with its length. In other words, we are storing a constant amount of information while the number of combinations of the state grows exponentially with the number of elements. The result is very promising because, to the best of our knowledge, no previous models with less than 1B model parameters have near-perfect accuracy at this length in this task.
6 Related Works
RNN-Based Language Models
This paper focuses on Mamba-2, a recurrent architecture that can be viewed as a variant of gated linear attention (Yang et al., 2024a). Many recently proposed RNNs can also be viewed as GLA variants. These include the RWKV series (Peng et al., 2023; 2024a), the Mamba series (Gu & Dao, 2023; Dao & Gu, 2024), GLA (Yang et al., 2024a), and many more (Zhang et al., 2024b; Yang et al., 2024b; De et al., 2024; Arora et al., 2024b; Orvieto et al., 2023; Sun et al., 2023). Our methods may apply to these architectures as well. Some recent/concurrent RNNs such as Gated DeltaNet (Yang et al., 2025), RWKV-7 (Peng et al., 2025), xLSTM (Beck et al., 2024), and Titans (Behrouz et al., 2024) have gone beyond a gating-based memory decay mechanism and are out of the scope of this paper.
Length Generalization
Most SOTA language models in the last few years have been based on the transformer (Vaswani et al., 2017) architecture. These models, when using certain variants of position encoding, can process arbitrarily long sequences. However, they exhibit severe performance drops on tokens beyond the training length (Zhao et al., 2024). To alleviate this shortcoming, many works have focused on modifying positional encoding (Peng et al., 2024b; Zhu et al., 2024; Ding et al., 2024; Jin et al., 2024), some achieving training-free length generalization to certain extents (Zhang et al., 2025).
Length Generalization of Mamba
Some prior works investigated the performance of Mamba as a function of context length (Park et al., 2024; Wen et al., 2025). Jelassi et al. (2024) empirically showed a sharp performance drop beyond the training length for Mamba on a copying task and also showed that Mamba struggles to copy from context unless its state size grows linearly with the context length. Arora et al. (2024a) discussed the associative recall abilities of transformer and some RNNs. Wang et al. (2025) is most related to our work. They discussed the issue of over-smoothing introduced by the memory decay term. In contrast, our paper explores a setting where recency may be preferred, but interference from earlier tokens damages the recall accuracy of recent tokens.
Some concurrent works have explored extending Mamba’s context length by controlling the discretization term ($\Delta_{t}$ in Eq. 2) (Ben-Kish et al., 2024), such as dividing it by a constant to make it smaller (Zhang, 2023). This makes the memory decay factor ($\alpha_{t}$ in Eq. 4) closer to 1, which makes the state retain more contextual information. However, it also unnecessarily diminishes the inserted information on all tokens. Similar to the above works, this study explores the cause of Mamba-2’s inability to generalize beyond its training context length and provides valuable insights into training Mamba-2 models that generalize better.
@@ 7 Conclusion
This paper demonstrates that while the Mamba architecture includes a memory decay mechanism, it fails to effectively learn forgetting in practice. As a result, when the context exceeds the training length, the model produces incoherent outputs. This issue arises from training with contexts that are too short relative to the state size. Empirical results show that robust forgetting is only learned when the training context length surpasses a certain threshold, which increases linearly with the state size. Notably, the model is still capable of recalling some contextual information beyond this threshold. These findings offer valuable insights into the causes and consequences of the model’s inability to forget, highlighting key limitations of the Mamba architecture. Nevertheless, the insights gained from this study provide a promising foundation for improving Mamba’s performance in long-context modeling, paving the way for more effective applications in tasks requiring extended context lengths.